a. The points 𝐴,𝐡 and 𝐢 have co-ordinates (βˆ’1,2),(1,1) and (2,3) respectively. Sketch the triangle 𝐴𝐡𝐢. By calculating the lengths of the sides of this triangle, determine if it is scalene, isosceles, or equilateral.


1
Expert's answer
2022-02-07T12:11:16-0500
"AB=\sqrt{(1-(-1))^2+(1-2)^2}=\sqrt{5}"

"BC=\sqrt{(2-1)^2+(3-1)^2}=\sqrt{5}"

"AC=\sqrt{(2-(-1))^2+(3-2)^2}=\sqrt{10}"

The triangle ABC is isosceles.